# NVS-Solver: Video Diffusion Model as Zero-Shot Novel View Synthesizer

### v1

@@ 4.3 NVS from Multiviews and Monocular Videos
We have illustrated the single view-based NVS-Solver in Algorithm 1, which can be further modified for NVS from multiple views or monocular videos. Specifically, given multiple views of a static scene and target poses, we first estimate the depth map of each of the given views and derive the warped view at target poses by warping the nearest given views to each target pose. For a monocular video, we warp each frame of the video sequence to the corresponding target pose with the same timestamp. Then, we sample the novel view via the proposed NVS-Solver as Lines 3-16 of Algorithm 1.
@@ 5 Experiment
Datasets. For single view-based NVS, we employed totally nine scenes, with six scenes from the Tanks and Temples dataset [22], containing both outdoor and indoor environments. The other three additional scenes are randomly chosen from the Internet. For multiview-based NVS, we used three scenes from the Tanks and Temples dataset [22], including both outdoor and indoor settings, as well as six scenes from the DTU dataset [15], which feature indoor objects. For each scene, we selected two images as input to perform view interpolation. For monocular video-based NVS, we downloaded nine videos from YouTube, each comprising 25 frames and capturing complex scenes in both urban and natural settings.
@@
Implementation Details. We conducted all the experiments with PyTorch using a single NVIDIA GeForce RTX A6000 GPU-48G. We adopted the point-based warping [40] to achieve $\mathcal{W}(\cdot)$ and employed Depth Anything [56] to estimate the maps of the input views. We simultaneously rendered 24 novel views and set the reverse steps as 100 for high-quality sample generation. For the implementation of Eq. (11), since applying directly weighted sum usually results in blurry, we ordered the feature pixels by the $\|\bm{\mu}_{t,\mathbf{p}_{i}}-\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}\|_{2}$ and take the ratio of $\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}$ smaller pixels from $\widehat{\mathbf{X}}_{0,\mathbf{p}_{0i}}$ and the others from $\bm{\mu}_{t,\mathbf{p}_{i}}$ to modulate $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$. We choose the values $(v_{1},v_{2},v_{3})$ as $(1e^{-6},9e^{-1},5e^{-2})$, $(1e^{-6},7e^{-1},1e^{-2})$, and $(1e^{-6},1.75,3e^{-2})$ for single, sparse, dynamic scene view synthesis.
@@
Metrics. We utilized four metrics to measure the reconstruction performance, i.e., Fréchet Inception Distance (FID) [12] evaluating the quality and diversity of synthesized views; Absolute trajectory error (ATE) [10] measuring the difference between the estimated trajectory of a camera or robot and the ground truth trajectory; Relative pose error (RPE) [10] measuring the drift, where we separately calculated the transition and rotation as RPE-T and RPE-R, respectively. Due to that the current depth estimation algorithms are hard to derive the absolute depth, we refrain from calculating image metric values like PSNR or SSIM.
@@
Table 1: Quantitative comparison of different methods on view synthesis, where we measure the FID and the error of view pose. We name our method with NVS-Solver (DGS) or (Post) for utilizing directly guided sampling in Eq. (13) or posterior sampling in Eq. (14), respectively. "*" indicates incomplete evaluation, where the corresponding metric cannot output a meaningful value on some sequences by the method due to the failure of pose estimation. “– –” denotes that the method cannot work on the condition or the metrics cannot be calculated. For all metrics, the lower, the better.
Methods Train Single view Multi-view
@@
Sparse Gaussian [55] 369.19 – – – – – – 324.60 – – – – – –
Sparse Nerf [49] – – – – – – – – 180.26 6.129 1.711 1.804
Text2Nerf [59] 187.05 2.223 0.718 0.107 – – – – – – – –
3D-aware [54] $\times$ 217.19 2.836 1.258 1.662 211.25 2.159 5.679 2.119
MotionCtrl [51] $\times$ 179.24 3.851 0.705 0.835 154.27 37.68 19.61 1.646
NVS-Solver (DGS) $\times$ 166.50 4.533 0.810 0.742 124.31 22.00 17.34 1.338
NVS-Solver (Post) $\times$ 165.12 0.767 0.156 0.170 126.44 4.052 2.030 0.330
Figure 2: Visual comparison of single view-based NVS results by (a) Text2Nerf [59], (b) 3D-aware [54], (c) MotionCtrl [51], (d) our NVS-Solver (Post). The middle view of each scene highlighted with the red rectangle refers to the input view. Here, we only show the results of the best two of all compared methods. We also refer reviewers to the Appendix and video demo contained in the supplementary file for more impressive results and comparisons.
Figure 3: (a) The two input views of each scene highlighted with the red rectangle. Visual results of multiview-based NVS by (b) 3D-aware [54], (c) MotionCtrl [51], (d) Our NVS-Solver (Post).
@@ 5.1 Results of NVS from Single or Multiple Views of Static Scenes
Fig. 2 visualizes synthesized novel views of two scenes by different methods from single views, where we can see that our NVS-Solver can consistently generate high-quality novel views with visually pleasing geometry and textures. The quantitative results listed in Table 1 also validate the significant superiority of our method over state-of-the-art methods. Although MontionCtrl [51] can generate high-quality images reflected by the FID value, its results have significantly large view pose errors reflected by the higher ATE and RPE values. For Text2Nerf [59], only 7 out of the total 9 sequences can be calculated metrics, which may be induced by the inherent errors of the learned geometric structure.
Fig. 3 shows the synthesized views by different methods from two given views, where it can be seen that our NVS-Solver outperforms state-of-the-art methods by clearer views, especially for the $2^{nd}$ and $3^{rd}$ scenes. In addition, it is worth noting that the lower ATE of 3D-aware [54] is due to the incomplete evaluation (See the supplementary material for more explanations).
Figure 4: Visual comparison on dynamic scene view synthesis of (a) input frames in the corresponding time of generated images, (b) Deformable-Gaussian [57], (c) 4D-Gaussian [52], (d) 3D-aware [54], (e) MotionCtrl [51], (f) Our NVS-Solver (Post).
@@ 5.2 Results of NVS from Monocular Videos of Dynamic Scenes
Deformable-Gaussian [57] 115.82 1.813 0.678 0.613
4D-Gaussian [52] 74.34 2.087 0.625 0.825
3D-aware [54] $\times$ 159.03 3.100 1.343 1.368
MotionCtrl [51] $\times$ 70.35 3.384 1.069 0.653
NVS-Solver (DGS) $\times$ 37.973 2.236 0.691 0.446
NVS-Solver (Post) $\times$ 39.86 2.308 0.725 0.400
The non-generative Gaussian-based methods Deformable-Gaussian [57] and 4D-Gaussian [52] usually cannot handle the marginal area, as illustrated in $2^{nd}$, $4^{th}$ and $6^{th}$ columns of Fig. 4 (b) and (c). Although the SVD-based method MotionCtrl can generate the boundary region as shown in Fig. 4 (e), the synthesized views are blurry and the poses of generated samples cannot follow the prompt, especially on the $3^{rd}$ sample. However, our method consistently generates high-quality novel views with lower pose errors, indicating its strong potential.
@@ 5.3 Ablation Study
Table 3: Quantitative comparison of different numbers of inference steps. All metrics are the lower the better.
@@
Posterior sampling vs. Directly guided sampling. The quantitative results listed in Tables 1 and 2 show that both the NVS-Solver (DGS) and NVS-Solver (Post) can generate high-quality images with comparable FID. However, the view pose of NVS-Solver (Post) is much more accurate than NVS-Solver (DGS), which is also visually verified by the results in Figs. 5 (c) and (d). Besides, NVS-Solver (DGS) takes 6 minutes for rendering 25 views, while NVS-Solver (Post) uses 1 hour.
Figure 5: Ablation studies of the proposed weight strategy of $\widehat{\lambda}(t,\mathbf{p}_{i})$ and embedding strategy. Visual comparison of (a) warped input image $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$, (b) results without our proposed weight strategy, (c) results of our NVS-Solver (DGS), and (d) results of our NVS-Solver (Post).
Figure 6: NVS results of our NVS-Solver (Post) with different trajectories, where the given view is bounded by a blue circle. We draw the pose at the right bottom corner of each sub-figure.
@@
Different Trajectories. We detailedly visualize the result of the proposed method on different trajectories. We apply transitions on eight different directions. The results are shown in Fig. 6. The proposed method can consistently render high-quality novel views.
@@ 6 Conclusion & Discussion
We have presented a training-free novel view synthesis paradigm, namely NVS-solver. NVS-Solver achieves remarkable performance compared with state-of-the-art methods. The advantages are credited to the powerful generative capacity of the pre-trained large stable video diffusion model and our elegant designs of the adaptive modulation of the diffusion score function through comprehensive theoretical and empirical analyses.
@@
Although our NVS-Solver takes longer than existing methods, the promising generative capacity inherent in the large pre-trained diffusion model may attract improvements in the future. Since the proposed method can synthesize views more accurate poses, we believe it may be a potential solution for the distillation of a pose controllable video diffusion model. In the age of generative intelligence, the authors expect the proposed algorithm would inspire future work to unify computer graphics pipelines with generative models. Finally, the authors sincerely appreciate Stability AI for the open-sourced SVD [5].
Appendix A Visual Comparison Details

### v2

@@ 4.3 NVS from Multiviews and Monocular Videos
We have illustrated the single view-based NVS via the proposed method in Algorithm 1, which can be further modified for NVS from multiple views or monocular videos. Specifically, given multiple views of a static scene and target poses, we first estimate the depth map of each of the given views and derive the warped view at target poses by warping the nearest given views to each target pose. For a monocular video, we warp each frame of the video sequence to the corresponding target pose with the same timestamp (See the Appendix E for the detailed warp strategy pipeline). Then, we sample the novel view via the proposed method as Lines 3-16 of Algorithm 1.
@@ 5 Experiment
Datasets. For single view-based NVS, we employed a total of nine scenes, with six scenes from the Tanks and Temples dataset (Knapitsch et al., 2017), containing both outdoor and indoor environments. The other three additional scenes are randomly chosen from the Internet. For multiview-based NVS, we used three scenes from the Tanks and Temples dataset (Knapitsch et al., 2017), including both outdoor and indoor settings, as well as six scenes from the DTU dataset (Jensen et al., 2014), which feature indoor objects. For each scene, we selected two images as input to perform view interpolation. For monocular video-based NVS, we downloaded nine videos from YouTube, each comprising 25 frames and capturing complex scenes in both urban and natural settings.
@@
Implementation Details. We conducted all the experiments with PyTorch using a single NVIDIA GeForce RTX A6000 GPU-48G. We adopted the point-based warping (Somraj, 2020) to achieve $\mathcal{W}(\cdot)$ and employed Depth Anything (yang2024depth) to estimate the maps of the input views (See the Appendix F for more results with different depth estimation methods). We simultaneously rendered 24 novel views and set the reverse steps as 100 for high-quality sample generation. For the implementation of Eq. (11), since applying directly weighted sum usually results in blurry, we ordered the feature pixels by the $\|\bm{\mu}_{t,\mathbf{p}_{i}}-\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}\|_{2}$ and take the ratio of $\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}$ smaller pixels from $\widehat{\mathbf{X}}_{0,\mathbf{p}_{0i}}$ and the others from $\bm{\mu}_{t,\mathbf{p}_{i}}$ to modulate $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$. We choose the values $(v_{1},v_{2},v_{3})$ as $(1e^{-6},9e^{-1},5e^{-2})$, $(1e^{-6},7e^{-1},1e^{-2})$, and $(1e^{-6},1.75,3e^{-2})$ for single, sparse, dynamic scene view synthesis.
@@
Metrics. We utilized four metrics to measure the reconstruction performance, i.e., Fréchet Inception Distance (FID) (Heusel et al., 2017) evaluating the quality and diversity of synthesized views; Absolute trajectory error (ATE) (Goel et al., 1999) measuring the difference between the estimated trajectory of a camera or robot and the ground truth trajectory; Relative pose error (RPE) (Goel et al., 1999) measuring the drift, where we separately calculated the transition and rotation as RPE-T and RPE-R, respectively. We utilize Particle-SFM (Zhao et al., 2022) to estimate the camera trajectory and assess the pose metrics. Since current depth estimation algorithms struggle to derive absolute depth from a single view or monocular video, resulting in a scale gap between the synthesized and ground truth images, we only compare the paired metrics in Appendix I, such as LPIPS.
@@
Table 1: Quantitative comparison of different methods on view synthesis, where we measure the FID and the error of view pose. We name our method with Ours (DGS) or (Post) for utilizing directly guided sampling in Eq. (13) or posterior sampling in Eq. (14), respectively. ”*” indicates incomplete evaluation, where the corresponding metric cannot output a meaningful value on some sequences by the method due to the failure of pose estimation. “– –” denotes that the method cannot work on the condition or the metrics cannot be calculated. For all metrics, the lower, the better.
Methods Overfitting Single view Multi-view
@@
Sparse Gaussian (Xiong et al., 2023) 369.19 – – – – – – 324.60 – – – – – –
Sparse Nerf (Wang et al., 2023a) – – – – – – – – 180.26 6.129 1.711 1.804
Text2Nerf (Zhang et al., 2024) 187.05 2.223 0.718 0.107 – – – – – – – –
Photoconsistent-NVS (Yu et al., 2023) $\times$ 193.87 7.64 1.19 1.45 – – – – – – – –
3D-aware (Xiang et al., 2023) $\times$ 217.19 2.836 1.258 1.662 211.25 2.159 5.679 2.119
MotionCtrl (Wang et al., 2024) $\times$ 179.24 3.851 0.705 0.835 154.27 37.68 19.61 1.646
Ours (DGS) $\times$ 166.50 4.533 0.810 0.742 124.31 22.00 17.34 1.338
Ours (Post) $\times$ 165.12 0.767 0.156 0.170 126.44 4.052 2.030 0.330
Figure 3: Visual comparison of single view-based NVS results by (a) Text2Nerf (Zhang et al., 2024), (b) 3D-aware (Xiang et al., 2023), (c) MotionCtrl (Wang et al., 2024), (d) Ours (Post). The middle view of each scene highlighted with the red rectangle refers to the input view. Here, we only show the results of the best two of all compared methods. We also refer reviewers to the Appendix B and video demo contained in the supplementary file for more impressive results and comparisons.
Figure 4: (a) The two input views of each scene highlighted with the red rectangle. Visual results of multiview-based NVS by (b) 3D-aware (Xiang et al., 2023), (c) MotionCtrl (Wang et al., 2024), (d) Ours (Post).
@@ 5.1 Results of NVS from Single or Multiple Views of Static Scenes
Fig. 3 visualizes synthesized novel views of two scenes by different methods from single views, where we can see that our method can consistently generate high-quality novel views with visually pleasing geometry and textures. The quantitative results listed in Table 1 also validate the significant superiority of our method over state-of-the-art methods. Although MontionCtrl (Wang et al., 2024) can generate high-quality images reflected by the FID value, its results have significantly large view pose errors reflected by the higher ATE and RPE values. For Text2Nerf (Zhang et al., 2024), only 7 out of the total 9 sequences can be calculated metrics, which may be induced by the inherent errors of the learned geometric structure.
Fig. 4 shows the synthesized views by different methods from two given views, where it can be seen that our method outperforms state-of-the-art methods by clearer views, especially for the $2^{nd}$ and $3^{rd}$ scenes. In addition, it is worth noting that the lower ATE of 3D-aware (Xiang et al., 2023) is due to the incomplete evaluation (See the Appendix C for more results).
Figure 5: The input views of each scene highlighted with the red rectangle. Visual results of synthesized 360° NVS from (a) single view and (b) multi-view input.
Figure 6: Visual comparison on dynamic scene view synthesis of (a) input frames in the corresponding time of generated images, (b) Deformable-Gaussian (Yang et al., 2024c), (c) 4D-Gaussian (Wu et al., 2024a), (d) 3D-aware (Xiang et al., 2023), (e) MotionCtrl (Wang et al., 2024), (f) Ours (Post).
Note that our method can also achieve 360° NVS through iterative execution of the proposed algorithm. Fig. 5 shows the synthesized 360° NVS from both single-view and multi-view inputs, demonstrating the capability of the proposed method to effectively handle the NVS task (See the Appendix H for the detailed strategy).
@@ 5.2 Results of NVS from Monocular Videos of Dynamic Scenes
Deformable-Gaussian (Yang et al., 2024c) 115.82 1.813 0.678 0.613
4D-Gaussian (Wu et al., 2024a) 74.34 2.087 0.625 0.825
3D-aware (Xiang et al., 2023) $\times$ 159.03 3.100 1.343 1.368
MotionCtrl (Wang et al., 2024) $\times$ 70.35 3.384 1.069 0.653
Ours (DGS) $\times$ 37.973 2.236 0.691 0.446
Ours (Post) $\times$ 39.86 2.308 0.725 0.400
The non-generative Gaussian-based methods Deformable-Gaussian (Yang et al., 2024c) and 4D-Gaussian (Wu et al., 2024a) usually cannot handle the marginal area, as illustrated in $2^{nd}$, $4^{th}$ and $6^{th}$ columns of Fig. 6 (b) and (c). Although the SVD-based method MotionCtrl can generate the boundary region as shown in Fig. 6 (e), the synthesized views are blurry and the poses of generated samples cannot follow the prompt, especially on the $3^{rd}$ sample. However, our method consistently generates high-quality novel views with lower pose errors, indicating its strong potential.
@@ 5.3 Ablation Study
Table 3: Quantitative comparison of different numbers of inference steps. For all metrics, the lower, the better.
@@
Posterior Sampling vs. Directly Guided Sampling. The quantitative results listed in Tables 1 and 2 show that both the Ours (DGS) and Ours (Post) can generate high-quality images with comparable FID. However, the view pose of Ours (Post) is much more accurate than Ours (DGS), which is also visually verified by the results in Figs. 7 (c) and (d). Besides, Ours (DGS) takes 6 minutes to render 25 views, while Ours (Post) uses 1 hour.
Table 4: Quantitative comparison of ablating updating rate $\kappa$ and the normalization schemes on Ours (Post). For all metrics, the lower, the better.
$\kappa$ Normalization FID ATE RPE-T RPE-R
$2e^{-2}\sqrt{\sigma(t)}$ Y 165.12 0.767 0.156 0.170
0.5 Y 268.17 1.92 0.318 0.399
$2e^{-2}\sqrt{\sigma(t)}$ N 176.00 4.71 0.568 1.05
Effectiveness of Back-propagation in Posterior Sampling. We performed ablation studies on the updating rate $\kappa$ and the normalization schemes in Eq. (14). The experimental results shown in Tab. 4 empirically guarantee that the back-propagation brings the latent space closer to the inherent low-dimensional manifold structure during each update step, further validating the performance of the proposed score-modulation schemes.
Figure 7: Ablation studies of the proposed weight strategy of $\widehat{\lambda}(t,\mathbf{p}_{i})$ and embedding strategy. Visual comparison of (a) warped input image $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$, (b) results without our proposed weight strategy, (c) results of Ours (DGS), and (d) results of Ours (Post).
Figure 8: NVS results of Ours (Post) with different trajectories, where the given view is bounded by a blue circle. We draw the pose at the right bottom corner of each sub-figure.
@@
Different Trajectories. We detailedly visualize the result of the proposed method on different trajectories. We apply transitions in eight different directions. The results are shown in Fig. 8. The proposed method can consistently render high-quality novel views.
@@ 6 Conclusion & Discussion
We have presented a training-free novel view synthesis paradigm. The proposed method achieves remarkable performance compared with state-of-the-art methods. The advantages are credited to the powerful generative capacity of the pre-trained large stable video diffusion model and our elegant designs of the adaptive modulation of the diffusion score function through comprehensive theoretical and empirical analyses.
@@
Although our method takes longer than existing methods, the promising generative capacity inherent in the large pre-trained diffusion model may attract improvements in the future. Since the proposed method can synthesize views more accurate poses, we believe it may be a potential solution for the distillation of a pose controllable video diffusion model. In the age of generative intelligence, the authors expect the proposed algorithm would inspire future work to unify computer graphics pipelines with generative models. Finally, the authors sincerely appreciate Stability AI for the open-sourced SVD (Blattmann et al., 2023).
Ethics & Reproducibility Statements
The proposed method is specifically designed for novel view synthesis through video diffusion. As such, no additional information regarding human subjects or potentially harmful insights is introduced during the process. This approach prioritizes privacy and ethical considerations by relying solely on the data available from the input images or videos, without requiring any sensitive or external information. Furthermore, all of our experiments are conducted in a training-free manner, which not only simplifies implementation but also ensures reproducibility. More importantly, the source code can be found on the webpage: https://github.com/ZHU-Zhiyu/NVS_Solver.
Appendix A Scene Prior-based Score-Modulation Preserves Data Manifold
A.1 Modulated Score Is More Accurate
In this section, we first illustrate that the proposed modulation is more accurate than guided sampling in an inpainting-like manner. Denote by $\widehat{\mathbf{X}}_{0,p_{i}}$ the warping content for scene’s prior on pose $p_{i}$. Then, the vanilla guided sampling methods, especially the inpainting-based method, directly utilize $\widehat{\mathbf{X}}_{0,p_{i}}$ as the guidance prior. Thus, its score estimation error $\mathcal{E}$ is shown as
$\mathcal{E}_{P}=\left\|\widehat{\bm{\mu}}_{t,\mathbf{p}_{i}}-\widehat{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\right\|_{2}=\left\|\widehat{\mathbf{X}}_{0,p_{i}}-\widehat{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\right\|_{2}=v_{3}\left\|\Delta\mathbf{p}_{i}\right\|_{2},$ (19)
where $\Delta\mathbf{p}_{i}$ indicates the variations of camera pose between the given and target views. However, for our modulated score, its score estimation error $\mathcal{E}_{M}$ is shown as
$\mathcal{E}_{M}$ $=\left\|\frac{1}{1+\lambda(t,\mathbf{p}_{i})}\bm{\mu}_{t,\mathbf{p}_{i}}+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}-\widehat{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\right\|_{2}$ (20)
$=\left\|\frac{1}{1+\lambda(t,\mathbf{p}_{i})}(\bm{\mu}_{t,\mathbf{p}_{i}}-\widehat{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*})+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}(\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}-\widehat{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*})\right\|_{2}$ (21)
$\leq\frac{1}{1+\lambda(t,\mathbf{p}_{i})}\left\|\bm{\mu}_{t,\mathbf{p}_{i}}-\widehat{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\right\|_{2}+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}\left\|\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}-\widehat{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\right\|_{2}$ (22)
$=\frac{v_{2}\sigma(t)}{1+\lambda(t,\mathbf{p}_{i})}+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}v_{3}\|\Delta\mathbf{p}\|_{2}$ (23)
By substituting the utilized solution of $\widehat{\lambda}(t,\mathbf{p}_{i})$ from Eq. (18), with $Q=v_{3}\|\Delta\mathbf{p}\|_{2}-v_{2}\sigma(t)$, we can derive that
$\mathcal{E}_{M}$ $\leq\frac{2v_{1}v_{2}\sigma(t)}{-Q+\sqrt{Q^{2}+4v_{1}Q}}+\frac{-(2v_{1}+Q)+\sqrt{Q^{2}+4v_{1}Q}}{-Q+\sqrt{Q^{2}+4v_{1}Q}}v_{3}\|\Delta\mathbf{p}\|_{2},$ (24)
$=\frac{2v_{1}v_{2}\sigma(t)}{-Q+\sqrt{Q^{2}+4v_{1}Q}}+\frac{-2v_{1}}{-Q+\sqrt{Q^{2}+4v_{1}Q}}v_{3}\|\Delta\mathbf{p}\|_{2}+\mathcal{E}_{P},$ (25)
$=\frac{-2v_{1}Q}{-Q+\sqrt{Q^{2}+4v_{1}Q}}+\mathcal{E}_{P}.$ (26)
since $v_{1}$ is a small value we make further approximation that
$\mathcal{E}_{P}$ $\leq\lim_{v_{1}\rightarrow 0}\frac{-2v_{1}Q}{-Q+\sqrt{Q^{2}+4v_{1}Q}}+\mathcal{E}_{P},$ (27)
${\approx}\lim_{v_{1}\rightarrow 0}\frac{-2v_{1}Q}{-Q+Q+\frac{1}{2\sqrt{Q^{2}+4v_{1}Q}}}+\mathcal{E}_{P},$ (28)
$=\lim_{v_{1}\rightarrow 0}\frac{-2v_{1}Q}{-Q+Q+\frac{1}{2\sqrt{Q^{2}+4v_{1}Q}}}+\mathcal{E}_{P},$ (29)
$=\lim_{v_{1}\rightarrow 0}-4v_{1}Q^{2}+\mathcal{E}_{P},$ (30)
$\approx\mathcal{E}_{P},$ (31)
where circled-1 indicates to apply Taylor series of $\sqrt{Q^{2}+4v_{1}Q}$.Thus, $\mathcal{E}_{M}\lesssim\mathcal{E}_{P}$ indicates that the proposed method modulates a more accurate target score function with less error.
A.2 Guided Posterior Sampling of Modulated Score Is on Data Manifold
According to Chung et al. (2023); Huang et al. (2022), we define a local tangent space as $\mathcal{T}_{\bm{x}}\mathcal{M}$ for a local orthogonal projection onto manifold $\mathcal{M}$, with a transition process
$\mathbf{Q}_{t}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},\bm{x}_{t}\rightarrow\bm{\mu}_{t}=\mathcal{X}_{\bm{\theta}}(\bm{x}_{t},t)$ (32)
$\frac{\partial}{\partial\bm{x}}\|\bm{\mu}_{t}-\widetilde{\bm{\mu}}_{t}\|_{2}^{2}=2\frac{\partial\bm{\mu}_{t}}{\partial\bm{x}}(\bm{\mu}_{t}-\widetilde{\bm{\mu}}_{t}).$ (33)
Since for the Jacobin matrix $\mathcal{J}_{\mathbf{Q}_{t}}=\frac{\partial\bm{\mu}_{t}}{\partial\bm{x}}$ denotes a transition which maps a vector to the tangent space of function $\mathbf{Q}_{t}$. Thus, we have
$\frac{\partial}{\partial\bm{x}}\|\bm{\mu}_{t}-\widetilde{\bm{\mu}}_{t}\|_{2}^{2}=\mathcal{J}_{\mathbf{Q}_{t}}\left(2(\bm{\mu}_{t}-\widetilde{\bm{\mu}}_{t})\right)\in T_{\mathbf{Q}_{i}}\mathcal{M},$ (34)
The aforementioned proof indicates that our introduced regularization, especially (DGS), only advocates moving the latent in the orthogonal manifold direction. Moreover, considering that the updating step size is relatively small as $2e^{-2}\sqrt{\sigma(t)}$ compared to the noised latent of $\sqrt{\sigma^{2}(t)+1}$, we have
$\frac{2e^{-2}\sqrt{\sigma(t)}}{\sqrt{\sigma^{2}(t)+1}}=\frac{2e^{-2}}{\sqrt{\sigma(t)+\frac{1}{\sigma(t)}}}\leq\frac{2e^{-2}}{\sqrt{2}}=\sqrt{2}e^{-2}.$ (35)
Thus, both the orthogonal regularization direction and small updating step make the proposed method to be safely moving on the data manifold.
Appendix B Visual Comparison Details